• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Benaoum/jpchen

    Log entry time 13:36:58 on July24,2005

    Entry number 149292

    keyword=Target Status

    Turned out the vent valve was not closed. So we closed the vent valve.
    Switched to a new bottle of helium before started cool-down.
    Target cooled-down completed.
    Normal condtions for target now : CHL flow around 20 (HP heater power=180w) and 200psi for pressure.
    The beam compensation turned back on. Fan turned back to 60 Hz.